# DCF-DS: Deep Cascade Fusion of Diarization and Separation for Speech Recognition under Realistic Single-Channel Conditions

### v1

@@ III-C Speaker Index-based Stitching
where and represent the estimated global T-F mask and the STFT of the mixed signal (including phase information) for the entire utterance, respectively. The index ranges from 1 to , where is the total frame number in the utterance as defined in Eq. (12). The separated time-domain signal is obtained by applying the inverse STFT to . After obtaining the separated signals for each speaker, re-clustering is performed to derive an enhanced global diarization prior:
This re-clustering process is illustrated on the right side of Fig. 2. One of the main advantages of re-clustering is that it reduces speaker confusion errors, particularly due to the reduced overlap regions in the separated speech, thereby improving the reliability of the speaker cluster. By leveraging these improved diarization priors, the overall system performance can be further enhanced.
@@ IV-A Datasets
Our primary focus is on real-world multi-speaker single-channel scenarios. Thus, we primarily evaluate the proposed method on the single-channel track of the NOTSOFAR-1 eval small dataset [6], consistent with the evaluation set used in the challenge, allowing direct comparison with other submitted systems. The NOTSOFAR-1 evaluation set comprises 80 real meeting recordings sampled at 16 kHz, each captured by two different single-channel devices. It consists of English conversations lasting approximately 6 minutes per session, involving 4 to 8 participants with balanced gender representation. The evaluation set reflects a wide range of real-world complexities, including diverse acoustic conditions such as varying speaker distances, fluctuating volumes, and multi-level of background noises, along with dynamic changes in the acoustic transfer function. It also presents conversational challenges like overlapping speech, rapid speaker shifts, interruptions, and non-verbal sounds. Collected across 10 rooms with distinct acoustic profiles, this dataset provides a robust basis for evaluating system performance in realistic environments. We use the development set in NOTSOFAR-1 datasets as our development set for hyper-parameter tuning.
For the training set, we first utilize the official 1000-hour NOTSOFAR-1 simulated dataset, which employs the Image Method [47] to simulate multi-channel data using clean speech from the Librivox corpus [48]. Since our focus is on training a single-channel system, we use the first channel for the training process. Additionally, we generate approximately 700 hours of simulated data using open-source scripts111https://github.com/jsalt2020-asrdiar/jsalt2020_simulate based on clean speech from the LibriSpeech corpus [49]. These two simulated datasets are used for joint training in DCF-DS, as well as for training the MIMO-SE module. Furthermore, the NOTSOFAR-1 challenge organizers have released a set of 53-hour real meeting training dataset. Although these real recordings lack clean source signals and cannot be directly used for training the separation model, we leverage them, in combination with the simulated training data, to train an independent NSD-MS2S model, which can provide more accurate diarization priors for the DCF-DS framework.
Additionally, to facilitate comparisons with more methods from the literature, we also evaluate our approach on the widely used LibriCSS corpus [18]. This dataset comprises 10 one-hour recordings of LibriSpeech utterances played back through loudspeakers, with overlap ratios ranging from 0 to 40 and involving 8 speakers per session. Consistent with previous studies [4, 43], we adopt the single-channel configuration in the publicly available script222https://github.com/chenzhuo1011/libri_css. It is worth noting that we test our trained model directly on the LibriCSS dataset without applying any specific domain adaptation.
@@ IV-B Implementation Details
We adopt the NOTSOFAR-1 challenge baseline system [6] as the baseline system on the NOTSOFAR-1 eval set, which follows the “CSS + SC” pipeline as described in Section II-B.
In the DCF-DS framework, we use a consistent window length of 64 ms and a hop size of 16 ms for the STFT transformation, resulting in a frequency dimension . The NSD-MS2S module processes 40-dimensional FBANK features, with an encoder composed of 6 Conformer blocks, each following the structure in [50]. The decoder module consists of 6 blocks with uniform settings: two attention layers with 512 dimensions and 8 attention heads, along with 1024-dimensional feed-forward layers. For the separation component, we adopt the Conformer architecture [43] from the NOTSOFAR-1 baseline [6] with one key modification: instead of using only the amplitude spectrum of the mixed speech as input, we include the amplitude spectrum and soft time masks. Our Conformer configuration features an attention dimension of 512, 8 attention heads, and 18 blocks.
During training, we use the Adam optimizer [51] with a learning rate of and train for 20 epochs (we also explore the performance with different numbers of epochs in preliminary experiments). The weight of the loss function in Eq. (24) is set to 1. We experiment with two window length configurations: seconds with a maximum of speakers per window, matching the NOTSOFAR-1 baseline for comparison, and seconds with a maximum of speakers, which allows the model to capture more context. To accelerate convergence, we initialize the diarization and separation modules in DCF-DS with pre-trained weights from NSD-MS2S and CSS, both trained on the simulated dataset. During the decoding phase, we use spectral clustering [52] to generate the global diarization prior .
For the LibriCSS dataset, we utilize the speaker embedding extractor from the ESPnet LibriCSS recipe [53]. For the NOSOFAR-1 dataset, speaker embeddings are extracted using a ResNet-221 model [54], which is trained on the VoxCeleb [55] and LibriSpeech datasets. Additionally, as mentioned in Section IV-A, we independently train an NSD-MS2S system on both simulated and real training datasets to provide more accurate priors, using the default configuration from the open-source code333https://github.com/liyunlongaaa/NSD-MS2S.
@@ IV-C Performance Measures
In our experiments, we primarily use the performance of the back-end ASR to evaluate the impacts of different front-end methods on speech quality. For the LibriCSS dataset, we evaluate ASR performance using the concatenated minimum-permutation word error rate (cpWER) [57]. cpWER is a widely used metric for multi-speaker scenarios which concatenates transcriptions for each speaker and finds the permutation that minimizes the WER. For the NOTSOFAR-1 eval dataset, we adopt the time-constrained minimum-permutation word error rate (tcpWER) [58] as our evaluation metric, consistent with the NOTSOFAR-1 challenge. TcpWER is an extension of cpWER that accounts for word errors, temporal boundary mismatches, and speaker identification errors. Even when words are correctly recognized, significant discrepancies in time boundaries result in penalties. Compared to cpWER, tcpWER incorporates temporal constraints and offers faster computation. In line with the official baseline in NOTSOFAR-1 challenge [6], we apply tcpWER with a 5-second collar. Additionally, to further analyze the proposed method, we measure the diarization error rate (DER) [59] for certain systems, which includes miss (MI), false alarm (FA), and confusion (CF) errors. We do not apply the forgiveness collars to the DER calculation.
@@ IV-D Performance Comparison on NOTSOFAR-1 Dataset
Table I shows the performance of different front-end systems with the same ASR back-end. For DCF-DS, we set the window length and the maximum speakers per window , consistent with the NOTSOFAR-1 baseline configuration. The training set for DCF-DS in this table is also restricted to the NOTSOFAR-1 simulated dataset, consistent with the baseline setup. From the table, we observe that using oracle diarization priors to segment the original mixed speech yields a TcpWER of 39.73. Even with oracle speaker boundaries, the segmented audios still contain substantial overlap regions, making it difficult to achieve high recognition accuracy and illustrating the dataset’s challenges. The NOTSOFAR-1 baseline system achieves a TcpWER of 41.4, slightly worse than the result obtained with oracle diarization priors. In the DCF-DS framework, using spectral clustering as the global prior results in a tcpWER of 38.02, representing an absolute improvement of 3.38 over the NOTSOFAR-1 baseline. We also compared performance across three different epochs (20, 30, and 40) and observed minimal differences among them. Furthermore, we incorporate an independently trained NSD-MS2S system, enhanced with real data, to provide the prior . This improved prior leads to consistent performance gains across epochs 20, 30, and 40, reducing the best TcpWER from 38.02 to 37.66. These results indicate that a more accurate diarization prior enhances the DCF-DS results, which is consistent with our expectations. Finally, we replace the outputs of the diarization module in DCF-DS, namely from Eq. (14), with the probabilities predicted by the independently trained NSD-MS2S system. This means substituting the “soft time mask” in Figures 2 and 3, which we refer to as “replace time mask (RTM)”. An interesting observation emerges: even though this independently trained NSD-MS2S model was not included in the joint training process, incorporating its predicted probabilities directly into the separation module of DCF-DS resulted in performance improvements. This can be attributed to the addition of real data in the training set, indicating that the DCF-DS framework can effectively leverage real-world data to enhance final results. Since we observe minimal performance differences across epochs 20, 30, and 40, we consistently use the results from epoch 20 in subsequent experiments for training efficiency. Moreover, among the different DCF-DS configurations, using RTM yields the best performance. Therefore, unless otherwise specified, all subsequent experiments use DCF-DS with RTM as the default configuration.
@@
Table II presents the performance of DCF-DS systems under different configurations. First, comparing the first and second blocks, we illustrate the impact of window lengths with the same NOTSOFAR-1 simulated training dataset. When we extend the window length from the baseline configuration of 3 seconds to 12.8 seconds (800 frames with a 16-ms frame shift), the maximum number of speakers per window () increases to 4. The results show that increasing the window length effectively improves system performance, reducing the tcpWER from 37.31 to 34.93. In the second block, we present results for a variant of the DCF-DS system where the time soft mask is replaced with i-vectors. This configuration is similar to the TS-SEP approach [4], where speaker embeddings are used to associate output with specific speaker identities, facilitating target-speaker separation. As shown in the results, relying solely on speaker embeddings without any temporal boundary information leads to a performance decline (tcpWER increasing from 34.93 to 37.35). This highlights the importance of the speaker diarization component in DCF-DS, as removing it significantly impacts performance, and underscores the positive role of time boundaries in enhancing the separated speech quality. Furthermore, [4] reports that directly training TS-SEP with an 8-output node structure results in non-convergence. In contrast, the 4-output node structure used in DCF-DS (i-vector) converges successfully, indicating that the proposed DCF-DS framework effectively mitigates the SDCI issue. In the third block, we first present results obtained by introducing additional simulated data based on the Librispeech dataset, referred to as “Libri-Sim”. Adding more simulated data yields only a slight improvement (tcpWER from 34.93 to 34.51), maybe because the original 1000-hour simulated data already provide sufficient coverage, so additional data do not introduce significant new information. After passing the separated streams through the MIMO-SE module, we observe an enhancement in speech quality, which effectively improves ASR performance (tcpWER from 34.51 to 33.81). However, when the T-F masks predicted by the first MIMO-SE module are fed iteratively back into MIMO-SE for a multi-stage process, as shown in Fig. 2, there is no additional gain (tcpWER remains at 33.97). This indicates that a single pass through MIMO-SE achieves sufficient speech quality enhancement, with no further improvement from additional iterations. Finally, we refine the initial speaker diarization priors through a re-clustering process (the details will be analyzed in the subsequent results). Initializing the DCF-DS system with these improved diarization priors leads to a substantial performance gain, reducing tcpWER from 33.97 to 31.72. Overall, our system achieves a relative improvement of 23.4 compared with the baseline, reducing tcpWER from 41.4 to 31.72.
@@
Fig. 4 presents the detailed DERs of different speaker diarization systems, along with the corresponding tcpWER results obtained from DCF-DS ( = 3s, = 3) without using RTM. Comparing the DERs of “SC” and “SC + NSD-MS2S”, we observe a significant improvement (from 27.62 to 21.09). This improvement mainly comes from reductions in MI (a decrease of 6.3, from 14.5 to 8.2). However, the tcpWER only shows a modest reduction of 0.37 (from 38.28 to 37.91). This suggests that while speaker diarization performance shows considerable improvement, it does not always lead to a significant enhancement in speech recognition performance, consistent with the conclusion in [14]. Moreover, comparing the “SC (Re-clustering)” and “SC + NSD-MS2S”, we observe that although the diarization result after re-clustering (DER = 23.81) is worse than the first NSD-MS2S result (DER = 21.09), the corresponding tcpWER shows a significant improvement (35.89 versus 37.91). This improvement can primarily be attributed to the reduction in CF (from 6.1 to 3.6), which nearly yields an equivalent enhancement in speech recognition performance (an absolute decrease of 2.02). Therefore, CF is a particularly important metric in speaker diarization results, especially for multi-speaker speech recognition. Furthermore, comparing the “SC (Re-clustering)” and “SC (Re-clustering) + NSD-MS2S”, we observe that although the overall DER of NSD-MS2S improves compared to the initialized SC system (decreasing from 23.81 to 19.72), its CF slightly increases (from 3.6 to 4.8), which is detrimental to the final speech recognition performance. This may be due to the fact that NSD-MS2S, compared to SC, reduces much of the MI, leading to the recovery of more speakers and inevitably increasing CF. From the perspective of speech recognition, initializing with NSD-MS2S still results in a slight improvement (tcpWER decreases from 35.89 to 35.79), likely because the significant reduction in MI by NSD-MS2S compensates for the increase in CF. Finally, comparing the “SC” and “SC (Re-clustering)”, we find that the use of re-clustering substantially improves diarization performance (DER reduced from 27.62 to 23.81), with the main gain coming from the decrease in CF (from 7.8 to 3.6). This demonstrates that the DCF-DS framework effectively leverages separated speech to enhance diarization performance, thereby bringing a significant improvement in downstream speech recognition.
@@
Table III presents the results of single systems submitted by other teams in the single-channel track of the NOTSOFAR-1 challenge. It is important to note that since each system may have used different training datasets and incorporated various improvements in the back-end systems, we are not making a strict performance comparison between methods. Instead, our goal is to evaluate whether the proposed approach can achieve performance on par with advanced methods in the field. In [60], the system used the EEND with transformer-based attractor calculation (EEND-TA) [65] method for speaker diarization and employed Whisper “large-v3” as the back-end, achieving a tcpWER of 44.2. In [61], the system utilized the PixIT framework [21] as the front-end and fine-tuned (FT) the Whisper “large-v3” on in-domain data for speech recognition, resulting in a tcpWER of 41.2. In [62], the system used the EEND system combined with WavLM as the front-end, and applied the Frame-Level Diarization Dependent Transformations (FDDT) method for the back-end Whisper model, transforming it into a target-speaker ASR. The ASR model was also fine-tuned on in-domain data, achieving a tcpWER of 40.1. Systems [63] and [64] both adopted the same “CSS + SC” approach as the baseline, with optimizations made to various modules. The system in [63] utilized the TF-GridNet [66] and multi-level normalized maximum eigengap-based (ML-NME) spectral clustering methods, and incorporated some ASR techniques such as rescoring using a pre-trained language model (LM), achieving a tcpWER of 33.5. The system in [64] combined WavLM with the CSS component in the front-end, and enhanced the speaker embedding extraction model with the ResNet293 [67]. The system in [64] also fine-tuned the back-end model, ultimately achieving a tcpWER of 34.6. The proposed DCF-DS system achieves the lowest tcpWER of 31.7 when using the baseline Whisper “large-v3” as the back-end recognition model. Furthermore, when the improved back-end model, Enhanced Whisper [54], is employed, the performance is further enhanced, achieving a tcpWER of 22.2. Ultimately, our system won first place in the single-channel track of the NOTSOFAR-1 Challenge666https://www.chimechallenge.org/challenges/chime8/task2/results.

### v2

@@ III-C Speaker Index-based Stitching
where and represent the estimated global T-F mask and the STFT of the mixed signal (including phase information) for the entire utterance, respectively. The frame index ranges from 1 to , where is the total frame number in the utterance as defined in Eq. (12). The separated time-domain signal is obtained by applying the inverse STFT to . After obtaining the separated signals for each speaker, re-clustering is performed to obtain an enhanced global diarization prior:
This re-clustering process is illustrated on the right side of Fig. 2. One of the main advantages of re-clustering is that it reduces speaker confusion errors. This is mainly due to the reduction of overlap regions in the separated speech, thus improving the reliability of the speaker clustering. By leveraging these improved diarization priors, the overall performance can be further enhanced.
@@ IV-A Datasets
Our primary focus is on real-world multi-speaker single-channel scenarios. Thus, we primarily evaluate the proposed method on the single-channel track of the NOTSOFAR-1 eval small dataset [6], consistent with the evaluation set used in the challenge, allowing direct comparison with other submitted systems. The NOTSOFAR-1 evaluation set comprises 80 real meeting recordings sampled at 16 kHz, each captured by two different single-channel devices. It consists of English conversations lasting approximately 6 minutes per session, involving 4 to 8 participants with balanced gender representation. The evaluation set covers a wide range of real-world complexities, such as diverse speaker distances, varying volumes, multi-level background noises, and dynamic changes in the acoustic transfer function. It also includes conversational challenges like overlapping speech, rapid speaker transitions, interruptions, and so on. Collected across 10 rooms with distinct acoustic profiles, this dataset provides a well-established benchmark for multi-speaker speech recognition in realistic environment. We use the development set in NOTSOFAR-1 dataset as our development set for hyper-parameter tuning. For the training set, we first utilize the official 1000-hour NOTSOFAR-1 simulated dataset, which employs the Image Method [47] to simulate multi-channel data using clean speech from the Librivox corpus [48]. Since our focus is on training a single-channel system, we use the first channel for the training process. Additionally, we generate approximately 700 hours of simulated data using open-source scripts111https://github.com/jsalt2020-asrdiar/jsalt2020_simulate based on the near-field close-talking recordings corpus in released training and development datasets. These two simulated datasets are used for joint training in DCF-DS, as well as for training the MIMO-SE module. Furthermore, the NOTSOFAR-1 challenge organizers have released a set of 53-hour real meeting training dataset. Although these real recordings lack clean source signals and cannot be directly used for training the separation model, we leverage them, in combination with the simulated training data, to train an independent NSD-MS2S model, which can provide more accurate diarization priors for the DCF-DS framework.
Additionally, to facilitate comparisons with more methods from the literature, we also evaluate our approach on the widely used LibriCSS corpus [18]. The libriCSS dataset comprises 10 one-hour recordings. These recordings contain many utterances from LibriSpeech, played back through loudspeakers in the meeting room, with overlap ratios ranging from 0 to 40. Each session involves 8 speakers. Consistent with previous studies [4, 43], we adopt the single-channel configuration in the publicly available script222https://github.com/chenzhuo1011/libri_css. It is worth noting that we test our trained model directly on the LibriCSS dataset without applying any specific domain adaptation.
@@ IV-B Implementation Details
We adopt the NOTSOFAR-1 challenge baseline system [6] as the baseline system for our evaluation on the NOTSOFAR-1 eval set, following the “CSS + SC” pipeline as described in Section II-B.
In the DCF-DS framework, we use a consistent window length of 64 ms and a hop size of 16 ms for the STFT transformation. The NSD-MS2S module processes 40-dimensional FBANK features, with the encoder composed of 6 Conformer blocks with the same structure as in [50]. The decoder module consists of 6 blocks with uniform settings: two attention layers with 512 dimensions and 8 attention heads, along with 1024-dimensional feed-forward layers. For the separation component, we adopt the Conformer architecture [43] from the NOTSOFAR-1 baseline [6] with one key modification: instead of using only the amplitude spectrum of the mixed speech as input, we include both the amplitude spectrum and soft time masks. Our Conformer configuration features an attention dimension of 512, 8 attention heads, and 18 blocks.
During training, we employ the Adam optimizer [51] with a learning rate of and conduct the training for 20 epochs (we also explore the performance with different numbers of epochs in preliminary experiments). The weight of the loss function in Eq. (24) is set to 1. We explore two different window length configurations: seconds with the maximum speaker number per window, aligning with the NOTSOFAR-1 baseline, and seconds with the maximum speaker number , which enables the model to capture more context. To accelerate convergence, we initialize the diarization and separation modules in DCF-DS with pre-trained weights from NSD-MS2S and CSS trained on the simulated datasets. During the decoding phase, we use spectral clustering [52] to generate the global diarization prior .
For the LibriCSS dataset, we utilize the speaker embedding extractor from the ESPnet LibriCSS recipe [53]. For the NOSOFAR-1 dataset, we use the overlap detection module to assist the SC in handling overlapping regions as in [54]. The speaker embeddings are extracted using a ResNet-221 model [54], which is trained on the VoxCeleb [55] and LibriSpeech datasets. Additionally, as mentioned in Section IV-A, we independently train an NSD-MS2S system on both simulated and real training datasets to provide more accurate priors, employing the default configuration from the open-source code333https://github.com/liyunlongaaa/NSD-MS2S.
@@ IV-C Performance Measures
In our experiments, we primarily evaluate the impacts of different front-end methods on speech quality using the back-end ASR performance. For the LibriCSS dataset, we evaluate ASR performance using the concatenated minimum-permutation word error rate (cpWER) [57]. cpWER is a widely used metric for multi-speaker scenarios which concatenates transcriptions for each speaker and finds the permutation that minimizes the WER. For the NOTSOFAR-1 eval dataset, we adopt the time-constrained minimum-permutation word error rate (tcpWER) [58] as our evaluation metric, consistent with the NOTSOFAR-1 challenge. The tcpWER is an extension of cpWER that accounts for word errors, temporal boundary mismatches, and speaker identification errors. Even when words are correctly recognized, significant discrepancies in time boundaries result in penalties. Compared to cpWER, tcpWER incorporates temporal constraints and offers faster computation. In line with the official baseline in NOTSOFAR-1 challenge [6], we apply tcpWER with a 5-second collar. Additionally, to further analyze the proposed method, we measure the diarization error rate (DER) [59] for certain systems, which includes miss (MI), false alarm (FA), and confusion (CF) errors. We do not apply the forgiveness collars to the DER calculation.
@@ IV-D Performance Comparison on NOTSOFAR-1 Dataset
Table I shows the performance of different front-end systems with the same ASR back-end. For DCF-DS, we set the window length and the maximum speakers per window , consistent with the NOTSOFAR-1 baseline configuration. The training set for DCF-DS in this table is also restricted to the NOTSOFAR-1 simulated dataset, consistent with the baseline setup. From the table, we can observe that utilizing oracle diarization priors to segment the original mixed speech yields a tcpWER of 39.73. Even with oracle speaker boundaries, the segmented audios still contain substantial overlap regions, making it difficult to achieve high recognition accuracy and illustrating the challenges of the dataset. The NOTSOFAR-1 baseline system achieves a tcpWER of 41.4, slightly worse than the result obtained with oracle diarization priors. In the DCF-DS framework, using spectral clustering as the global prior results in a tcpWER of 38.02, representing an absolute improvement of 3.38 over the NOTSOFAR-1 baseline. We also compared performance across three different epochs (20, 30, and 40) and observed minimal differences among them. Furthermore, we incorporate an independently trained NSD-MS2S system, enhanced with real data, to provide the prior . This improved prior leads to consistent performance gains across epochs 20, 30, and 40, reducing the best tcpWER from 38.02 to 37.66. These results indicate that a more accurate diarization prior enhances the DCF-DS results, which is consistent with our expectations. Finally, we replace the outputs of the diarization module in DCF-DS, namely from Eq. (14), with the probabilities predicted by the independently trained NSD-MS2S system. This means substituting the “soft time mask” in Figures 2 and 3, which we refer to as “replace time mask (RTM)”. An interesting observation emerges: even though this independently trained NSD-MS2S model was not included in the joint training process, incorporating its predicted probabilities directly into the separation module of DCF-DS resulted in performance improvements. This can be attributed to the addition of real data in the training set, indicating that the DCF-DS framework can effectively leverage real-world data to enhance final results. Since we observe minimal performance differences across epochs 20, 30, and 40, we consistently use the results from epoch 20 in subsequent experiments for training efficiency. Moreover, among the different DCF-DS configurations, using RTM yields the best performance. Therefore, unless otherwise specified, all subsequent experiments use DCF-DS with RTM as the default configuration.
@@
Table II presents the performance of DCF-DS systems under different configurations. First, comparing the first and second blocks, we illustrate the impact of window length with the same NOTSOFAR-1 simulated training dataset. When we extend the window length from the baseline configuration of 3 seconds to 12.8 seconds (800 frames with a 16-ms frame shift), the maximum number of speakers per window () increases to 4. The results show that increasing the window length effectively improves system performance, reducing the tcpWER from 37.31 to 34.93. In the second block, we present results for a variant of the DCF-DS system where the time soft mask is replaced with the i-vectors. This configuration is similar to the TS-SEP approach [4], where speaker embeddings are used to associate outputs with specific speaker identities, facilitating target-speaker separation. As shown in the results, relying solely on speaker embeddings without any temporal boundary information leads to a performance decline (tcpWER increasing from 34.93 to 37.35). This highlights the importance of the speaker diarization component in DCF-DS, as removing it significantly impacts performance, and underscores the positive role of time boundaries in enhancing the separated speech quality. Furthermore, [4] reports that directly training TS-SEP with an 8-output node structure results in non-convergence. In contrast, the 4-output node structure used in DCF-DS (i-vector) converges successfully, indicating that the proposed DCF-DS framework effectively mitigates the SDCI issue. In the third block, we first present results obtained by introducing additional training data simulated from the near-field (NF) close-talking recordings corpus, referred to as “NF-Sim”. Adding more simulated data yields only a slight improvement (tcpWER from 34.93 to 34.51), maybe because the original 1000-hour simulated data already provide sufficient coverage, so additional data do not introduce significant new information. After passing the separated streams through the MIMO-SE module, we observe an enhancement in speech quality, which effectively improves ASR performance (tcpWER from 34.51 to 33.81). However, when the T-F masks predicted by the first MIMO-SE module are fed iteratively back into MIMO-SE for a multi-stage process, as shown in Fig. 2, there is no additional gain (tcpWER remains at 33.97). This indicates that a single pass through MIMO-SE achieves sufficient speech quality enhancement, with no further improvement from additional iterations. Finally, we refine the initial speaker diarization priors through a re-clustering process (the details will be analyzed in the subsequent results). Initializing the DCF-DS system with these improved diarization priors leads to a substantial performance gain, reducing tcpWER from 33.81 to 31.72. Overall, our system achieves a relative improvement of 23.4 compared with the baseline, reducing tcpWER from 41.4 to 31.72.
@@
Fig. 4 presents the detailed DERs of different speaker diarization systems, along with the corresponding tcpWER results obtained from DCF-DS ( = 3s, = 3) without using RTM. Comparing the DERs of “SC” and “SC + NSD-MS2S”, we observe a significant improvement (from 27.62 to 21.09). This improvement mainly comes from reductions in MI (a decrease of 6.3, from 14.5 to 8.2). However, the tcpWER only shows a modest reduction of 0.37 (from 38.28 to 37.91). This suggests that while speaker diarization performance shows considerable improvement, it does not always lead to a significant enhancement in speech recognition performance, consistent with the conclusion in [14]. Moreover, comparing the “SC (Re-clustering)” and “SC + NSD-MS2S”, we observe that although the diarization result after re-clustering (DER = 23.81) is worse than the first NSD-MS2S result (DER = 21.09), the corresponding tcpWER shows a significant improvement (35.89 versus 37.91). This improvement can primarily be attributed to the reduction in CF (from 6.1 to 3.6), which nearly yields an equivalent gain in speech recognition performance (an absolute decrease of 2.02). Therefore, CF is a particularly important metric in speaker diarization results, especially for multi-speaker speech recognition. Furthermore, comparing the “SC (Re-clustering)” and “SC (Re-clustering) + NSD-MS2S”, we observe that although the overall DER of NSD-MS2S improves compared to the initialized SC system (decreasing from 23.81 to 19.72), its CF slightly increases (from 3.6 to 4.8), which is harmful to the final speech recognition performance. This may be due to the fact that NSD-MS2S, compared to SC, reduces much of the MI, leading to the recovery of more speakers and inevitably increasing CF. From the perspective of speech recognition, initializing with NSD-MS2S still results in a slight improvement (tcpWER decreases from 35.89 to 35.79), likely because the significant reduction in MI by NSD-MS2S compensates for the increase in CF. Finally, comparing the “SC” and “SC (Re-clustering)”, we find that the use of re-clustering substantially improves diarization performance (DER reduced from 27.62 to 23.81), with the main gain coming from the decrease in CF (from 7.8 to 3.6). This demonstrates that the DCF-DS framework effectively leverages separated speech to enhance diarization performance, thereby bringing a significant improvement in downstream speech recognition.
@@
Table III presents the results of single systems submitted by other teams in the single-channel track of the NOTSOFAR-1 challenge. It is important to note that since each system may have used different training datasets and incorporated various improvements for the back-end systems, we are not making a strict performance comparison between methods. Instead, our goal is to present the performance obtained by advanced methods in the field. In [60], the system used the EEND with transformer-based attractor calculation (EEND-TA) [66] method for speaker diarization and employed Whisper “large-v3” as the back-end, achieving a tcpWER of 44.2. In [61], the system utilized the PixIT framework [21] as the front-end and fine-tuned (FT) the Whisper “large-v3” on in-domain data for speech recognition, resulting in a tcpWER of 41.2. In [62], the system used the EEND system combined with WavLM as the front-end, and applied the Frame-Level Diarization Dependent Transformations (FDDT) method for the back-end Whisper model, transforming it into a target-speaker ASR. The ASR model was also fine-tuned on in-domain data, achieving a tcpWER of 40.1. Systems [63], [64] and [65] all adopted the same “CSS + SC” approach as the baseline, with optimizations made to various modules. The system in [63] adopted the Zipformer transducer as the ASR model with WavLM-large features, and used the Conformer CSS and normalized maximum eigengap-based (NME) spectral clustering, achieving a tcpWER of 36.6. The system in [64] utilized the TF-GridNet [67] and multi-level normalized maximum eigengap-based (ML-NME) spectral clustering methods, and incorporated some ASR techniques such as rescoring using a pre-trained language model (LM), achieving a tcpWER of 33.5. The system in [65] combined WavLM with the CSS component in the front-end, and enhanced the speaker embedding extraction model with the ResNet293 [68]. The system in [65] also fine-tuned the back-end model, ultimately achieving a tcpWER of 34.6. The proposed DCF-DS system achieves the lowest tcpWER of 31.7 when using the baseline Whisper “large-v3” as the back-end ASR model. Furthermore, when the improved back-end model, Enhanced Whisper [54], is employed, the performance is further enhanced, achieving a tcpWER of 22.2. Ultimately, our system won first place in the single-channel track of the NOTSOFAR-1 Challenge666https://www.chimechallenge.org/challenges/chime8/task2/results.